OpenStudy (anonymous):

Give an example of an even function and explain algebraically why it is even.

OpenStudy (anonymous):

Can you guys help me explain it algebraically?

OpenStudy (anonymous):

Anyone?

OpenStudy (zzr0ck3r):

\(f(x) = x^2\) is even because \(f(-x) = (-x)^2=x^2=f(x)\)

OpenStudy (anonymous):

Thanks!